OpenStudy (anonymous):

if f(x) = { 2+x , x>=0 2-x , x<0 Find f(f(x)) ?

OpenStudy (anonymous):

@mahmit2012 @JFraser @akash123

OpenStudy (anonymous):

f(x)=2+abs(x) ff(x)=2+abs(f(x))=2+f(x)=4+abs(x)

OpenStudy (anonymous):

f>0 so absf=f

OpenStudy (anonymous):

f(f(x) will have two answer one is when x >=0 and x <0

terenzreignz (terenzreignz):

Not quite, see, if x is less than 0, then f(x) turns out to always be greater than zero. Don't take my word for it, though :D Try it :D

OpenStudy (anonymous):

4+x x>=0 4-x x<0 !!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

hw 4 -x x<0 ???

OpenStudy (anonymous):

the answer is 4+abs(x)

OpenStudy (anonymous):

@Yahoo!!!!!

OpenStudy (anonymous):

Yup....) thx....